Table 1 presents data on the opinion of a group of adults about astrology. There are two variables; "Opinion" with 2 categories (Not-scientific and Sort-of-scientific), and "Degree" with 3 categories (Ju- nior, Bachelor and Graduate). The numbers in the table are the counts of adults in each level of Opinion by each level of Degree, however, there are some missing counts Table 1 Degree Opinion Junior Bachelor | Graduate Total Not scientific 122 268 Very or sort of scientific 106 31 13 1. If "Opinion" and "Degree" are independent, what are the values of the missing counts? (explain your answer) 2. Find the marginal distribution of Degree
Actually, the Two-way table of observed counts is the following. Table 2: Degree Opinion Junior Bachelor| Graduate Not scientific 122 44 102 Very or sort of scientific 31 62 13 3. Find the marginal distributions of Opinion 4. Find the conditional distributions of Opinion, given each value of Degree (write your answer in a table) 5. Is Opinion associated to Degree? (explain your answer) What do you conclude?
Answer #1

1.

Junior Bachelor Graduate Row Total
Not scientific 78 122 68 268
scientific 31 62 13 106
Column Total 109 184 81 374

If Opinion and degree are independent then

cell = (Row total)*(column total)/Total

that is , 31 = (106)*(column total)/ 374

therefore column total = 109 (rounding)

2. Marginal distribution of degree

Degree Probability
Junior 109/374= 0.29
Bachelor 184/374=0.49
Graduate 81/374=0.22

3.

Junior Bachelor Graduate Row Total
Not scientific 44 122 102 268
scientific 31 62 13 106
Column Total 75 184 115 374

Marginal distribution of Opinion

Opinion Probability
Not scientific 268/374=0.72
Scientific 106/374=0.28

4.

Junior Bachelor Graduate
Not Scientific 44/75=0.59 122/184=0.66 102/115=0.89
Scientific 31/75=0.41 62/184=0.34 13/115= 0.11

Note : P(Not scientific I Junior ) = P(Not scientific \bigcap Junior)/ P(Junior)= (44/374)/(75/374) = 0.59

5.We got that

P(Not scientific I Junior )= 0.59

P(Not scientific) = 0.72

Thus ,

P(Not scientific I Junior ) \neq P(Not scientific)

We can say that Opinion is not independent of Degree

We can conclude that , Opinion is associated with degree